# Q-PETR: Quant-aware Position Embedding Transformation for Multi-View 3D Object Detection

### v1

@@ Abstract
PETR-based methods have dominated benchmarks in 3D perception and are increasingly becoming a key component in modern autonomous driving systems. However, their quantization performance significantly degrades when INT8 inference is required, with a degradation of 58.2% in mAP and 36.9% in NDS on the NuScenes dataset. To address this issue, we propose a quantization-aware position embedding transformation for multi-view 3D object detection, termed Q-PETR. Q-PETR offers a quantization-friendly and deployment-friendly architecture while preserving the original performance of PETR. It substantially narrows the accuracy gap between INT8 and FP32 inference for PETR-series methods. Without bells and whistles, our approach reduces the mAP and NDS drop to within 1% under standard 8-bit per-tensor post-training quantization. Furthermore, our method exceeds the performance of the original PETR in terms of floating-point precision. Extensive experiments across a variety of PETR-series models demonstrate its broad generalization.
@@ 1 Introduction
3D perception is crucial for autonomous driving and embodied intelligence [17, 67, 31, 16, 62]. PETR-based methods [49, 29, 30, 43, 56, 50, 52] have gained significant attention due to their effectiveness in this domain. Unlike previous detection methods relying on dense convolutional features [17, 61, 60, 59, 37] or utilizing 3D-to-2D projections as in DETR3D-based approaches [51, 25, 28], PETR adapts the paradigm of 2D object detection with transformers (DETR) [5] by incorporating 3D positional encoding, achieving seamless end-to-end 3D obstacle detection.
Despite its advantages, PETR requires substantial storage and computational resources during inference, posing challenges for on-device deployment. To enhance inference efficiency, considerable efforts have been dedicated to efficient model compression [42, 20, 27, 19, 11, 12], with quantization being particularly favored for deploying models on AI chips by representing networks in low-bit formats [39, 32, 54, 7].
However, deploying complex models like PETR on edge devices presents additional challenges. Autonomous vehicles rely on edge AI chips with limited computational capabilities. Nonlinear functions such as softmax, sigmoid, SiLU, and GeLU are commonly used during inference, but most edge AI chips lack sufficient floating-point performance for these operations, exacerbating deployment difficulties.
Moreover, quantization awareness is often overlooked in neural architecture design, leading to significant performance degradation during deployment. Notably, models like MobileNet [40, 64, 15, 38], EfficientNet [4, 13, 45], and RepVGG [8] face quantization issues requiring additional design efforts or advanced strategies to mitigate deployment challenges. PETR is no exception; when deployed on in-vehicle AI chips, standard 8-bit per-tensor post-training quantization (PTQ) results in a significant performance drop, with a 58.2% decrease in mAP and a 36.9% decrease in NDS.
In this paper, we address the quantization issues of PETR to facilitate its deployment on edge AI chips in autonomous vehicles. We conduct an in-depth analysis of the mechanisms behind PETR’s quantization failure, revealing that the significantly larger magnitudes of positional encodings compared to image features, and the highly imbalanced scaled dot-products in cross-attention, are the primary causes of quantization collapse. Based on these insights, we propose Q-PETR (Quantization-aware PETR), a model that mitigates quantization collapse and even improves floating-point performance compared to the original PETR.
Our contributions are summarized as follows:
Identifying the root causes of quantization collapse in PETR: We uncover that the imbalanced output of the inverse-sigmoid function and excessively large values in positional encoding, as well as the imbalanced scaled dot-products in cross-attention, are key issues.
Redesigning the positional encoding and quantization strategy: We reformulate the positional encoding and improve the quantization approach for the scaled dot-product in cross-attention, enhancing both floating-point performance and quantization friendliness.
Introducing DuLUT for efficient nonlinear function inference: We propose DuLUT, an improved lookup table [46] approach that reduces the number of table entries without compromising precision, optimizing nonlinear function inference on edge AI chips.
Demonstrating generalizability and deployment readiness: Our method generalizes well across different model scales, achieving exceptional post-quantization performance suitable for deployment on edge devices.
@@ 2 Related Work
3D Object Detection. Surround-view 3D object detection is essential for autonomous driving and is generally categorized into LSS-based [17, 59, 16] and transformer-based [29, 43] approaches. LSS-based methods project multi-camera features onto dense BEV (Bird’s Eye View) representations [37], but their high memory consumption hinders efficient long-range perception. Transformer-based methods leverage sparsity to enhance long-distance perception. Among these, the PETR series has gained significant attention. PETR [29] transforms 2D image features into 3D representations using 3D positional encoding. PETRv2 [30] introduces temporal feature indexing, while StreamPETR [49] extends temporal query processing. Some works [48, 52, 9] accelerate processing by incorporating 2D detection priors. CMT [56] fuses vision and LiDAR point clouds. Improvements to PETR’s positional encoding have also been explored [43, 14]. Additionally, PETR has been integrated into the Omnidrive framework [50] to enhance 3D perception with large models.
@@
with and being the maximum and minimum floating-point values from the calibration dataset. Quantization methods are categorized into Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT [10, 4] introduces quantization-aware losses during training, enhancing robustness but requiring resource-intensive retraining. PTQ offers rapid deployment without retraining. While PTQ methods have been successful on CNNs [33, 34, 21], they often perform poorly on Vision Transformers (ViTs) and transformer-based detectors due to structural differences. For ViTs, practical PTQ algorithms have been developed [63, 26, 44, 24, 41]. In the context of transformer-based object detection models, Q-DETR [55] and AQ-DETR [47] use QAT and knowledge distillation to mitigate performance degradation in low-bit quantization of DETR models. These methods primarily focus on quantizing GEMM operations. For nonlinear activation functions, lookup table (LUT) techniques [46] are commonly used. Additionally, methods like I-BERT [18] and I-ViT [22] employ integer approximation to achieve fixed-point computation.
Quantization for 3D Object Detection. Quantization methods have been applied to accelerate 3D object detection in autonomous driving and robotics. Leveraging advances in image quantization, QD-BEV [66] employs QAT and distillation in multi-camera 3D detection, achieving smaller models and faster inference than the BEVFormer baseline [23]. For LiDAR-based detection, LIDAR-PTQ [68] achieves state-of-the-art quantization on CenterPoint [58], with performance close to FP32 and a threefold speedup. To our knowledge, no PTQ solution currently exists for transformer-based 3D object detection in autonomous driving.
@@ 3 Quantization and Deployment-Friendly Adaptation of PETR
In this section, we aim to improve PETR’s quantization performance. We begin by introducing the principles of PETR (§3.1), identify its quantization failures (§3.2), and provide strategies to address these challenges (§3.3).
3.1 Preliminaries
PETR
@@ 3.2 Quantization Failure of PETR
Here, is the number of calibration data points; denotes the full-precision network; the quantization error is ; and represents the network output with quantization applied at bit-width , while other layers remain at full precision.
Since 8-bit weight quantization incurs minimal information loss, we focus on quantization errors related to operator inputs. Using the PETR configuration from the first row of Table 1, we obtain layer-wise SQNRs, depicted in Fig. 1. From these results, we identify three primary observations contributing to quantization errors:
Observation 1: The inverse-sigmoid operator disrupts feature distribution balance.
As indicated by the red arrow in Fig.1, quantization difficulties stem from PETR’s positional encoding module. Analyzing its construction (Fig.5(a)), we find that the inverse-sigmoid operation induces an imbalanced feature distribution. Specifically, Fig. 2 shows that before applying inverse-sigmoid, the feature distribution is balanced and quantization-friendly, whereas afterward, it exhibits significant outliers. Thus, avoiding the inverse-sigmoid operator in the positional encoder is advisable.
Observation 2: Disparity in magnitude between camera-ray PE and image features leads to information loss after quantization.
The issue highlighted by the purple arrows in Fig.1 indicates that applying 8-bit symmetric linear quantization to the 3D position-aware key significantly degrades performance. Since is derived by element-wise summing image features and camera-ray PE, we quantize them using ’s quantization parameters, as visualized in Fig.3. Clearly, image features occupy only 4 bits, causing severe information loss. Designing a positional encoding with a magnitude distribution similar to image features is crucial to address this quantization issue.
Observation 3: The scaled dot-product in cross-attention exhibits significant fluctuations along the head dimension.
As highlighted by the green arrow in Fig.1, the scaled dot-product in cross-attention presents quantization challenges. Since the subsequent softmax operation is applied along the last dimension, we visualize the maximum values over this dimension in Fig.4. The feature distribution varies significantly across different heads, posing challenges for effective per-tensor quantization.
@@
Based on the analysis in Section 3.2, we identify two key areas for improvement: (1) designing a positional encoding with magnitude and distribution similar to image features, and (2) balancing the scaled dot-products in cross-attention across heads. We propose specific methods to address these challenges.
@@
To reduce the magnitude of positional encoding in PETR, we analyze its construction process (Fig.5(a)). PETR samples 64 3D points along each camera ray per pixel, increasing magnitude variance. According to statistical theory, the variance of a sum increases with the number of variables. Inspired by 3DPPE [43], which introduces a lidar-centered lidar-ray PE using only one 3D point (Fig.5(b)), we design a quantization and deployment-friendly (QD-aware) lidar-ray PE (Fig.5(c)).
Our QD-aware lidar-ray PE avoids both the inverse-sigmoid operator and sinusoidal calculations, making it suitable for quantization and deployment on edge devices. Specifically, we learn three anchor embeddings along each axis , denoted with anchor locations for . For each lidar-ray point , we identify the two closest anchors and compute interpolated embeddings:
The QD-aware lidar-ray PE is then obtained via , where denotes concatenation along the channel dimension, and the MLP has a conv-relu-conv structure.
@@
In softmax operations, numerical stabilization (NS) subtracts the maximum value to prevent overflow. Traditional quantization quantizes before NS, leading to issues in Observation3. We propose quantizing after NS (Fig.6), and adaptively determining the optimal truncation lower bound to minimize softmax error.
After NS, inputs to softmax are non-positive. Values below approach zero after exponentiation, so we define a candidate set of scaling factors with for -bit quantization. The dequantized input is:
@@
Lookup table (LUT) is commonly used for integer inference of non-linear functions in quantized neural networks. Linear LUT can compute functions losslessly when the input quantization bit width is less than or equal to the number of LUT entries. However, for higher-bit inputs, conventional linear LUT become impractical due to exponentially increasing table sizes—requiring 256 entries for 8-bit inputs and 65,536 entries for 16-bit inputs—leading to significant storage and computational overhead.
To improve precision without excessive resource usage, non-linear LUT have been proposed. These methods allocate more entries to regions where the function changes rapidly and fewer entries where it remains flat, thus enhancing quantization precision in critical areas. However, non-linear LUT are complex to implement in hardware, necessitating additional processing to define input regions based on function characteristics, and are not efficiently supported by most hardware architectures. Approximation methods like I-BERT [18] and I-ViT [22] attempt to simplify computations but introduce approximation errors and increased computational steps, offsetting the benefits of quantization.
@@
To address these challenges, we propose DuLUT, which uses two small linear tables tailored to the function’s properties, enabling accurate non-linear approximation with significantly reduced table sizes. DuLUT splits the input domain into two regions—typically where the function is rapidly changing and where it is relatively flat—and employs a separate small linear LUT for each region. For example, with 8-bit quantization, DuLUT uses two tables of 32 entries each without compromising precision (see Fig.7 and Algorithm 2). We applied DuLUT to common activation functions like softmax, GELU, and SiLU. By utilizing DuLUT, we achieve the same precision as larger single-table lookups while significantly reducing SRAM overhead and maintaining computational efficiency.

### v2

# Q-PETR: Quantization-aware Position Embedding Transformation for Multi-View 3D Object Detection

@@ Abstract
Camera-based multi-view 3D detection has emerged as an attractive solution for autonomous driving due to its low cost and broad applicability. However, despite the strong performance of PETR-based methods in 3D perception benchmarks, their direct INT8 quantization for onboard deployment leads to drastic accuracy drops—up to 58.2% in mAP and 36.9% in NDS on the NuScenes dataset. In this work, we propose Q-PETR, a quantization-aware position embedding transformation that re-engineers key components of the PETR framework to reconcile the discrepancy between the dynamic ranges of positional encodings and image features, and to adapt the cross-attention mechanism for low-bit inference. By redesigning the positional encoding module and introducing an adaptive quantization strategy, Q-PETR maintains floating-point performance with a performance degradation of less than 1% under standard 8-bit per-tensor post-training quantization. Moreover, compared to its FP32 counterpart, Q-PETR achieves a two-fold speedup and reduces memory usage by three times, thereby offering a deployment-friendly solution for resource-constrained onboard devices. Extensive experiments across various PETR-series models validate the strong generalization and practical benefits of our approach.
@@ 1 Introduction
3D object detection [58, 59, 61, 26] has been a longstanding topic in computer vision. Compared to LiDAR, cameras have gained increasing popularity in autonomous systems due to their ability to provide dense texture information at a lower cost, making camera-based 3D object detection more favored [13, 12, 54]. Among these studies, the mainstream PETR frameworks [39, 23, 24, 34, 47, 40, 42] have gained prominence by adapting the 2D transformer-based DETR paradigm [5] with 3D positional encodings. In comparison to dense feature methods [13, 53, 52, 50, 32] or DETR3D-style 3D-to-2D projections [41, 19, 21], PETR achieves end-to-end 3D detection while performing superior performance. Despite its effective, the deployment of PETR on resource-limited edge devices presents a critical challenge that significantly hinders their widespread application in autonomous vehicles and robotics.
Quantization [60, 57, 7, 29] is an efficient model compression approach that reduces computational burden by converting high-bit floating-point into low-bit integer formats. Compared to quantization-aware training (QAT) methods, which require access to all labeled training data and substantial computation resources, post-training quantization (PTQ) is more suitable for rapid and efficient paratical applications. This is because PTQ only needs a small number of unlabeled samples for calibration. Furthermore, PTQ does not require retraining the network with all available labeled dataset, resulting in a shorter implementation time. Although several advanced PTQ methods have been proposed for 2D detection tasks[15, 43, 22] and ViT [16, 25], directly applying them to multi-view 3D Detection tasks inevitably leads to severe performance degradation due to the structures and task-specific differences. For instance, when standard 8-bit per-tensor post-training quantization (PTQ) is applied to PETR, leading up to 58.2% mAP and 36.9% NDS performance collapse.
Furthremore, nonlinear operators such as softmax, gelu, and silu are indispensable in 3D detection models, but their performance is often hindered by hardware constraints, akin to the “short board” in a barrel. Even specialized Tensor Cores in high-end GPUs like the NVIDIA A100 exhibit significantly lower throughput for these nonlinear operators compared to matrix multiplications [9]. Moreover, many edge AI chips rely on a lookup table (LUT) [36] for integer activation functions, but these typically only support integer inputs and are often linear LUTs. While a linear LUT can faithfully approximate nonlinearities if its size covers the entire dynamic range, its capacity grows exponentially (e.g., from 256 entries at 8-bit to 65536 entries at 16-bit), making it impractical. Non-linear LUTs [51] allocate more entries to steeper regions of the function and fewer to flatter regions, but introduce additional hardware complexity. Other integer-only methods [14, 16] further simplify operator emulation at the cost of increased approximation error. Consequently, effective quantization and hardware-friendly acceleration of these nonlinear operators is crucial and unexplored for PETR in resource-constrained deployment.
In this paper, we address these quantization challenges for PETR-based 3D Detection. Firstly, through an in-depth analysis, we find that disproportionately large positional encodings and imbalanced scaled dot-products in cross-attention significantly degrade quantized performance. Building on these findings, we propose Q-PETR, a quantization-friendly variant of PETR that not only mitigates performance collapse but also enhances floating-point accuracy. To further resolve the bottleneck of nonlinear operators, we introduce a lightweight dual-LUT (DuLUT) mechanism that maintains high approximation fidelity with fewer table entries. Our main contributions are summarized as follows:
Diagnosis of quantization failures in PETR: We show that large positional encodings, imbalanced inverse-sigmoid outputs, and skewed cross-attention dot-products are key factors causing significant accuracy loss under low-bit quantization.
Redesign of positional encodings and cross-attention quantization: We reformulate the positional encoding module and employ a more balanced scaling strategy for cross-attention dot-products, improving both floating-point and quantized performance.
Introduction of DuLUT for hardware-friendly nonlinear functions: By splitting LUT entries based on the curvature of the function, DuLUT efficiently approximates nonlinearities with fewer table entries, greatly facilitating edge deployment.
Broad applicability and deployment readiness: Our approach generalizes to various model scales, achieving minimal performance loss under standard 8-bit PTQ while even boosting full-precision accuracy, thus meeting the demands of resource-limited scenarios in real-world.
@@ 2 Related Work
Multi-View 3D Object Detection. Surround-view 3D object detection is essential for autonomous driving and is generally categorized into LSS-based [13, 50, 12] and transformer-based [23, 34] approaches. LSS-based methods project multi-camera features onto dense BEV (Bird’s Eye View) representations [32], but their high memory consumption hinders efficient long-range perception. Transformer-based methods leverage sparsity to enhance long-distance perception. Among these, the PETR series has gained significant attention. PETR [23] transforms 2D image features into 3D representations using 3D positional encoding. PETRv2 [24] introduces temporal feature indexing, while StreamPETR [39] extends temporal query processing. Some works [38, 42, 8] accelerate processing by incorporating 2D detection priors. CMT [47] fuses vision and LiDAR point clouds. Improvements to PETR’s positional encoding have also been explored [34, 11]. Additionally, PETR has been integrated into the Omnidrive framework [40] to enhance 3D perception with large models.
@@
with and being the maximum and minimum floating-point values from the calibration dataset. Quantization methods are categorized into Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT [10, 4] introduces quantization-aware losses during training, enhancing robustness but requiring resource-intensive retraining. Compare to QAT, PTQ offers rapid deployment without retraining. While PTQ methods have been successful on CNNs [28, 29, 15], they often perform poorly on transformer-based 3D detectors due to structural differences. For ViTs, practical PTQ algorithms have been developed [55, 20, 35, 18]. In the context of transformer-based object detection models, Q-DETR [46] and AQ-DETR [37] use QAT and knowledge distillation to mitigate performance degradation in low-bit quantization of DETR models. These methods primarily focus on quantizing GEMM operations. For nonlinear activation functions, lookup table (LUT) techniques [36] are commonly used. Additionally, methods like I-BERT [14] and I-ViT [16] employ integer approximation to achieve fixed-point computation.
Quantization for 3D Object Detection. Quantization methods have been applied to accelerate 3D object detection in autonomous driving and robotics. Leveraging advances in image quantization, QD-BEV [57] employs QAT and distillation in multi-camera 3D detection, achieving smaller models and faster inference than the BEVFormer baseline [17]. For LiDAR-based detection, LIDAR-PTQ [60] achieves state-of-the-art quantization on CenterPoint [49], with performance close to FP32 and 3 speedup. To our knowledge, there are no PTQ solution tailored for transformer-based 3D detection in autonomous driving.
@@ 3 Quantization and Deployment-Friendly Adaptation of PETR
In this section, we aim to improve PETR’s quantization performance. We begin by elaborating the principles of PETR in Sup. A, identify its quantization failures (§3.1), and provide strategies to address these challenges (§3.2).
@@ 3.1 Quantization Failure of PETR
Here, is the number of calibration data points; denotes the full-precision network; the quantization error is ; and denotes the network output when only the target layer is quantized to bits, with all other layers kept at full precision.
Since 8-bit weight quantization results in only a minor loss of precision, we focus on quantization errors arising from operator inputs. Using the PETR configuration from the first row of Table 1, we obtain layer-wise SQNRs, depicted in Fig. 1. From these results, we identify three main factors contributing to quantization errors:
Observation 1: Position Encoding Design Flaws Lead to Quantization Difficulties.
We find that PETR’s quantization issues primarily arise from its positional encoding module in two key ways. (a) Inverse-sigmoid disrupts feature balance. As shown in Fig. 1, the inverse-sigmoid operation skews an otherwise balanced distribution (Fig. 2) toward significant outliers. (b) Magnitude disparity between camera-ray PE and image features. As highlighted by the purple arrow in Fig. 1, applying 8-bit quantization to the 3D position-aware key yields severe performance drops. Statistical analysis (Fig. 3) shows that camera-ray PE spans approximately , while image features remain within . Consequently, when using Eq. 2 and an 8-bit symmetric range of , PE dominates the scaling factor. Image features then collapse into merely seven bins (Fig. 4), causing catastrophic information loss and sharp accuracy degradation (Table 1). To address these flaws, we (1) remove the inverse-sigmoid step that drives outlier magnitudes, and (2) redesign the positional encoding to align its scale with that of image features. This balanced approach preserves critical information during quantization.
Observation 2: Dual-Dimensional Heterogeneity in Cross-Attention Leads to Quantization Bottlenecks.
As evidenced by the green arrow in Fig. 1 and further clarified in Fig. 5, the scaled dot-product in cross-attention exhibits pronounced heterogeneity on two levels. First, the inter-head variance spans 2–3 orders of magnitude, while within each head, the value distribution is extremely broad (e.g., ranging beyond [, ]). We merge the head and query dimensions to directly reveal the row-wise feature distribution. The results show that regardless of whether quantization is performed per head, per token, or on the entire tensor, the excessively large softmax inputs result in significant quantization errors. This confirms that existing quantization paradigms are fundamentally inadequate for handling the severe amplitude disparities in the cross-attention mechanism.
@@
Drawing on the analysis in Section 3.1 and the deployment challenges noted in the Introduction, we identify three critical issues. Firstly, the positional encoding module mismatches the magnitude and distribution of image features, causing severe quantization loss. Secondly, imbalanced scaled dot-products in cross-attention further compound quantization errors. Thirdly, the high computational cost of nonlinear functions impedes efficient edge inference. We propose targeted solutions for above challenges.
@@
From the derivations in Appendix C.2, we establish that the amplitude of camera-ray PE can theoretically reach up to 11.5 times that of its LiDAR-based counterpart. This stark discrepancy directly explains why PETR’s camera-ray encoding often overshoots the dynamic range of image features, thereby hampering quantization. Although LiDAR-ray PE alleviates the amplitude issue, its reliance on high-frequency sinusoidal functions remains problematic for low-bit deployments on edge devices.
To overcome both amplitude and implementation obstacles, we propose a quantization-deployable LiDAR-ray position embedding (QDPE) that sharply curtails magnitudes while avoiding complex nonlinearities. Our design contains two main modifications:
1.
(Single-point sampling via LiDAR prior). Drawing inspiration from the physical properties of LiDAR sensors, we sample only one 3D point per pixel along each depth ray (Fig. 6 (b)), in contrast to the multi-sample scheme in PETR’s camera-ray PE. By discarding the iterative inverse-sigmoid and sinusoidal transformations, we reduce the overall encoding variance.
2.
(Anchor-based bounded embedding with convex-combination constraints). As illustrated in Fig. 6 (c), we learn three axis-aligned anchor embeddings for each spatial axis , with corresponding anchor locations . For a LiDAR-sampled 3D point , we compute the embedding along each axis by linear interpolation between the nearest two anchors:
(4) Theorem 1 guarantees that each axis-wise component is strictly confined within the convex hull of its adjacent anchors. We concatenate the three axis-wise embeddings and feed them to a lightweight MLP to obtain the final positional encoding vector.
These two innovations ensure our QD-aware LiDAR-ray PE remains both bounded in magnitude and free of difficult-to-quantize nonlinearities. Fig. 3 and Fig. 9 visually demonstrate the elimination of outliers. Further, the dynamic range of our QD-aware encoding () is only marginally wider than that of standard image features ()—in stark contrast to PETR’s original (). The proposed design eliminates complex nonlinear operations (inverse-sigmoid) and spectral components (high-frequency sinusoids), achieving hardware-compatible computation without compromising geometric fidelity.
@@
In softmax operations, numerical stabilization (NS) subtracts the maximum value to prevent overflow. Traditional quantization quantizes before NS, leading to issues in Observation2. We propose quantizing after NS (Fig. 7), and adaptively determining the optimal truncation lower bound to minimize softmax error.
After NS, inputs for softmax are non-positive. Values below approach zero after exponentiation, so we define a candidate set of scaling factors with for -bit quantization. The dequantized input is:
@@
The error bound of linear LUT can be formally expressed by the maximum interpolation error theorem. Given a twice-differentiable function over interval , the maximum approximation error using linear interpolation satisfies:
where represents the curvature. This result indicates that if the second derivative (i.e., curvature) is large within a sub-interval, must be shortened to control the approximation error. Conversely, if the curvature is small, the sub-interval can be lengthened. Consequently, more interpolation points (LUT entries) should be assigned where the function changes rapidly, while flatter or near-saturated regions may be merged into fewer intervals.
Building on this principle, DuLUT partitions the input domain into three types of sub-intervals—shrink (near-saturated), enlarge (steep change), and unchange (near-linear)—defined as follows:
shrink: for regions where the function is close to saturation or changes very little, multiple quantization steps are “compressed” into a single or few LUT entries;
enlarge: for high-curvature regions, more LUT entries are assigned to preserve accuracy;
unchange: for intervals that appear nearly linear, further subdivision is unnecessary.
As a result, extra entries can be concentrated in critical intervals (e.g., [-9, 8] for SiLU) to capture rapid nonlinear variations, while intervals far from the main dynamic range (e.g., , where the function is saturated) are merged. Let the high-curvature region have length enlarge_length, the full input domain be , and the total number of LUT entries be table_n. In a single linear LUT scheme, the number of entries assigned to the high-curvature region follows:
For example, if the SiLU function spans (hence ) and we employ a 512-entry linear LUT, this formula indicates that only entries would fall within the high-curvature region.
By contrast, DuLUT retains the same total of 512 entries but splits them into two smaller 256-entry LUTs. The first LUT maps the input into a “nonlinear index,” while the second LUT stores the actual function values. Continuing the SiLU example, if we allocate 256 entries following the above principle, the enlarge region might occupy 4 entries, the shrink region 126 entries (with only 1 used explicitly), and thus the enlarge region ultimately gains 129 entries. In effect, this yields a lookup resolution equivalent to using approximately 7588 entries in a single-table design.
@@
A example, with 8-bit quantization, DuLUT uses two tables of 32 entries each without compromising precision (see Fig. 8 and Algorithm 1). We applied DuLUT to common activation functions like softmax, GELU, and SiLU. By utilizing DuLUT, we achieve the same precision as larger single-table lookups while significantly reducing SRAM overhead and maintaining computational efficiency.
| Bac | Size | Feat | Mode | mAP | NDS | mATE | mASE | mAOE | mAVE | mAAE | |
| R50 | 512 1408 | c5 | PETR | FP32 | 31.42 | 36.11 | 84.19 | 28.42 | 60.69 | 99.08 | 23.58 |
| PTQ | 13.79 | 25.31 | 107.94 | 31.47 | 75.22 | 83.71 | 25.45 | ||||
| Q- PETR | FP32 | 31.49 | 37.20 | 82.52 | 27.88 | 59.91 | 91.74 | 23.45 | |||
| PTQ | 31.34 | 37.17 | 82.61 | 27.93 | 60.00 | 91.79 | 23.45 | ||||
| R50 | 512 1408 | p4 | PETR | FP32 | 32.60 | 37.16 | 82.63 | 27.96 | 61.06 | 95.81 | 23.91 |
| PTQ | 12.97 | 24.75 | 108.28 | 31.76 | 79.57 | 78.90 | 27.14 | ||||
| Q- PETR | FP32 | 32.69 | 38.03 | 80.58 | 27.89 | 59.43 | 92.69 | 22.55 | |||
| PTQ | 32.40 | 37.72 | 81.11 | 27.92 | 60.02 | 92.76 | 22.59 | ||||
| R101 | 512 1408 | p4 | PETR | FP32 | 34.40 | 38.62 | 80.67 | 28.03 | 57.13 | 95.74 | 24.20 |
| PTQ | 13.53 | 23.84 | 111.04 | 31.27 | 78.94 | 92.92 | 26.14 | ||||
| Q- PETR | FP32 | 34.72 | 39.68 | 79.40 | 27.92 | 53.90 | 92.99 | 22.59 | |||
| PTQ | 34.41 | 39.08 | 80.98 | 28.00 | 54.41 | 92.62 | 22.70 | ||||
| V2-99 | 320 800 | p4 | PETR | FP32 | 38.01 | 42.56 | 75.79 | 26.84 | 50.58 | 90.13 | 21.07 |
| PTQ | 10.46 | 23.64 | 112.41 | 33.00 | 85.96 | 71.83 | 25.12 | ||||
| Q- PETR | FP32 | 38.43 | 43.80 | 74.79 | 27.29 | 49.76 | 82.11 | 20.15 | |||
| PTQ | 37.93 | 43.17 | 75.50 | 27.78 | 50.07 | 82.41 | 20.38 | ||||
| V2-99 | 640 1600 | p4 | PETR | FP32 | 40.66 | 46.05 | 71.76 | 27.07 | 42.23 | 80.68 | 21.06 |
| PTQ | 6.40 | 20.98 | 117.38 | 34.85 | 83.38 | 76.83 | 27.10 | ||||
| Q- PETR | FP32 | 41.35 | 45.99 | 72.18 | 26.91 | 45.05 | 82.03 | 20.67 | |||
| PTQ | 40.95 | 45.64 | 73.40 | 27.05 | 45.61 | 82.17 | 20.68 | ||||
| V2-99 | 320 800 | p4 | StreamPETR | FP32 | 48.19 | 57.11 | 60.99 | 25.58 | 37.54 | 26.28 | 19.43 |
| PTQ | 18.52 | 36.47 | 76.39 | 31.44 | 47.03 | 30.22 | 20.99 | ||||
| Q- StreamPETR | FP32 | 49.13 | 57.57 | 60.77 | 26.14 | 39.05 | 24.81 | 19.15 | |||
| PTQ | 48.21 | 56.33 | 63.00 | 26.35 | 39.17 | 24.90 | 19.19 | ||||
| V2-99 | 640 1600 | p4 | StreamPETR | FP32 | 49.51 | 58.03 | 60.10 | 26.07 | 35.65 | 25.91 | 19.60 |
| PTQ | 18.72 | 35.66 | 74.32 | 30.39 | 41.49 | 30.53 | 20.82 | ||||
| Q- StreamPETR | FP32 | 50.48 | 58.61 | 58.78 | 26.16 | 37.05 | 25.69 | 18.59 | |||
| PTQ | 49.44 | 57.94 | 60.30 | 26.55 | 37.07 | 25.87 | 18.59 | ||||